HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

HIGH-SPEED VIBRATIONAL SPECTROSCOPY WITH TIME-STRETCH TECHNOLOGY

\underline{T. IDEGUCHI} , Institute for Photon Science and Technology, The University of Tokyo, 113-0033 Tokyo, Japan

The advancement of manipulation technologies for ultrashort laser pulses has led to significant improvements in the measurement speed of broadband vibrational spectroscopy. Laser frequency comb spectroscopy, particularly dual-comb spectroscopy, has dramatically increased the scan rate of interferometric measurements, thereby enabling high-speed Fourier-transform spectroscopy. However, its measurement speed is limited to approximately 1 MSpectra/s due to the signal-to-noise ratio. We have recently pioneered even faster vibrational spectroscopy techniques, capable of running at tens of MSpectra/s, by applying time-stretch technology. In this talk, I will introduce our latest developments in both infrared and Raman spectroscopy: timestretch infrared spectroscopy (TS-IR) and time-stretch coherent Raman spectroscopy (TS-CRS).

p-number: p235  Submitted on Thu, 22 Jun 2023 02:55:53 +0200